# FedAIoT: A Federated Learning Benchmark for Artificial Intelligence of Things

### v2

@@ A.3 Dataset Details
A.3.1 WISDM
The WISDM dataset comprises raw accelerometer and gyroscope data collected from 51 subjects performing 18 activities for three minutes each. Data was gathered at a 20Hz sampling rate from both a smartphone (Google Nexus 5/5x or Samsung Galaxy S5) and a smartwatch (LG G Watch). Data for each device and sensor type are stored in different directories, resulting in four directories overall. Each directory contains 51 files, each corresponding to a subject. The data entry format is: <subject-id, activity-code, timestamp, x, y, z>. Separate files for the gyroscope and accelerometer readings are provided and are later combined by matching timestamps. Subject ID is given from 1600 to 1650 and the activity code is an alphabetical character between ‘A’ and ‘S’ excluding ‘N’. The timestamp is in Unix time. The code to read and partition the data into 10s segments is provided by our benchmark. The input shape of the processed data is $200\times 6$. The original dataset is available at https://archive.ics.uci.edu/dataset/507/wisdm+smartphone+and+smartwatch+activity+and+biometrics+dataset.
A.3.2 UT-HAR
A.3.3 Widar
A.3.4 VisDrone
A.3.5 CASAS
The CASAS dataset is a collection of data generated in smart home environments, where intelligent software uses sensors deployed at homes to monitor resident activities and conditions within the space. The CASAS project considers environments as intelligent agents and employs custom IoT hardware known as Smart Home in a Box (SHiB), which encompasses the necessary sensors, devices, and software. The sensors in SHiB perceive the status of residents and their surroundings, and through controllers, the system acts to enhance living conditions by optimizing comfort, safety, and productivity. The CASAS dataset includes the date (in yyyy-mm-dd format), time (in hh:mm:ss.ms format), sensor name, sensor readings, and an activity label in string format. The data was collected in real-time as residents go about their daily activities. The code to extract categorical sensor readings to create input sequences and labels is provided in our benchmark. The CASAS dataset can be downloaded from https://casas.wsu.edu/datasets/.
A.3.6 AEP
A.3.7 EPIC-SOUNDS
A.4 Hyperparameters
@@
A.5 Model Architectures
A.5.1 WISDM
A.5.2 UT-HAR
A.5.3 Widar
A.5.4 VisDrone
A.5.5 CASAS
A.5.6 AEP
A.5.7 EPIC-SOUNDS

### v3

@@ B.7 EPIC-SOUNDS
Appendix C Formal Definition of Dirichlet Distribution
where $\Gamma(\cdot)$ is the gamma function.
Appendix D Model Architectures
We evaluated multiple model architectures for each dataset and selected the one that performed the best. Table 11 lists the models examined for each dataset with the selected model marked in bold.
D.1 WISDM
D.2 UT-HAR
D.3 Widar
D.4 VisDrone
D.5 CASAS
D.6 AEP
D.7 EPIC-SOUNDS
Table 11: List of models examined for each dataset. Selected model marked in bold.
{tblr}
width = colspec = Q[133]Q[158]Q[158]Q[120]Q[138]Q[142]Q[117], cells = c, hlines, WISDM & UT-HAR WIDAR VisDrone Casas AEP Epic Sounds
LSTM,
GRU
3-layer 1D CNN 3-layer Dense,
LeNet,
Resnet-18,
LSTM,
GRU,
BiLSTM,
3 layer 1D CNN,
2 layer Transformer 3-layer Dense,
LeNet,
Resnet-18,
LSTM,
GRU,
BiLSTM,
3 layer 1D CNN,
2 layer Transformer Yolo v3,
Yolo v5,
Yolo v8 LSTM,
BiLSTM,
3-layer LSTM,
3-layer BiLSTM 5-layer MLP,
SVM,
Linear Regressor ResNet18,
VGG-16
Appendix E Hyperparameters